Erin's age is 3 times Warren's. In 4 years she will be twice as old as he will be. How old is each now?
Answer by EMStelley(208)

Let's call Warren's age w. Then Emily is 3w years old. In 4 years from now, Warren will be w+4 and Emily will be 3w+4. If Emily will be twice as old as Warren is in 4 years, we have that
3w+4=2(w+4)
3w+4=2w+8
w=4
So Warren is 4 and Emily is 12.